The sum of 10 terms of the series whose nth terms is n-5. Please help.

Grade:12

1 Answers

P S S C ASLESH
2 years ago
nth term of the series is n²-5
Sum of n terms=∑(n²-5)=∑n²-5
=n(n+1)(2n+1)/6 -5
n=10
Sum of 10 terms = 10(10+1)((2*10)+1)/6 -5
=(10*11*21)/6 -5
=2310/6 -5
=385-5
=380